To practice Problem-Solving Strategy for general problems.

Two hockey pucks, labeled A and B, are initially at rest on a smooth ice surface and are separated by a distance of 18.0 m . Simultaneously, each puck is given a quick push, and they begin to slide directly toward each other. Puck A moves with a speed of 1.50 m/s , and puck B moves with a speed of 1.90 m/s .

What is the distance covered by puck A by the time the two pucks collide?

What is i, the distance that puck A covers prior to the collision? Express your answer in meters to three significant figures. Do not include units in your answer. View Available Hint(s) 16.6 SubmitPra Previous Answers X Incorrect; Try Again; 7 attempts remaining
Answer #1

Given : vA = 1.50 m/s , vB = 1.90 m/s , d = 18 m

Solution:

Since , two hockey pucks A and B are moving towards each other , their combined or relative velocity is given by :

V = VA+VB

= 1.50 m/s + 1.90 m/s

= 3.40 m/s

The time required to collide , is given by:

t = d/V

= (18 m)÷(3.40 m/s)

= 5.29 s

The distance that puck A covers prior to collision is given by:

x1 = VA×t

= (1.50 m/s)×(5.29 s)

= 7.94 m

Answer : x1 = 7.94 m